### v2

@@ 1 Introduction
Empirical evaluations conducted on the CrossDocked2020 dataset [36] demonstrate that MolPIF has comprehensive advantages across four key dimensions: (1) Superior de novo generation capability, producing molecules with superior binding properties and chemical validity; (2) Accurate geometric reproduction of molecular structural distributions, including rings, bond lengths, and bond angles; (3) Flexible adaptation to different prior distributions, where we systematically compared and analyzed the advantages and limitations of using Gaussian versus Laplace distributions as priors for modeling atomic coordinates; (4) Effective lead optimization, demonstrating robust performance in enhancing drug candidate properties.
@@ 2.4.3 Case analysis for the performance of MolPIF in de novo molecule generation
For the case study, we selected three representative binding pockets (2V3R, 1L3L, and 6VO5) for de novo molecule generation. Three-dimensional visualization of the generation results is presented in Figure 2. Figure 2a–d illustrate the generation outcomes for pockets 2V3R and 1L3L. The displayed molecules represented the top-performing candidates from 100 generated samples per model, selected according to a comprehensive ranking that considered all evaluation metrics. Quantitative analysis demonstrated that MolPIF-generated molecules exhibited superior performance across multiple metrics under the constraint of a fixed number of atoms. This improved ligand efficiency suggested substantial advantages for downstream optimization and experimental synthesis, substantiating the practical applicability and potential of MolPIF in real-world drug discovery applications.
To evaluate the model’s performance in a biologically relevant context, we applied MolPIF to generate potential inhibitors targeting HAT1. HAT1 is an enzyme that catalyzes the acetylation, which plays a critical role in various physiological processes, and its dysregulation has been closely associated with multiple human diseases, particularly cancers [55]. The generated molecules were then compared with both the reference molecule from the crystal structure and the preliminary screening hits reported in the literature. In this evaluation, MolPIF generated 100 candidate molecules targeting the HAT1 binding site (PDB ID: 6VO5). Remarkably, approximately 50% of these molecules outperformed molecule H9 in key metrics (Figure 2e,f), with the vast majority showing improvement over the reference. Notably, H9 was identified from 100,000 PocketFlow-generated molecules [56] and has been experimentally validated for its biological activity ($IC_{\text{50}}=72.36\mu M$). These results indirectly demonstrated the exceptional efficiency of MolPIF in de novo generation of active molecules as starting points for novel targets.
@@ 2.4.5 The performance of MolPIF in lead optimization
MolPIF demonstrated applicability not only to de novo molecular generation tasks but also to atom-level lead optimization. This capability enabled arbitrary specification of fixed atoms within given molecules while allowing the model to generate desired substituent regions. Figure 4 presents selected lead optimization cases performed on reference molecules targeting 1umd, 3ZCW, and 6KZZ.
The target 1umd was selected from the CrossDocked test set, with fixed ligand atom selection following the protocol established by CBGBench [54]. Figure 4a displays some molecular structures generated by MolPIF for pocket 1umd during lead optimization. The reference ligand was partitioned into distinct regions for four subtasks: fragment growth, linker design, scaffold hopping, and side-chain decoration. MolPIF exhibited dual functional capabilities: (1) extending existing scaffold groups through Fragment and Side Chain modifications, and (2) integrating discrete fragments into complete molecular structures via Linker and Scaffold operations. Notably, the model accomplished molecular optimization without explicit gradient guidance from property prediction, relying solely on learned structural information. This suggested that the model implicitly captured structural features associated with favorable molecular properties. Such atom-level lead optimization facilitated straightforward modification or replacement of molecular substructures, enabling optimization of physicochemical properties or circumvention of patent restrictions, thereby demonstrating significant practical potential.
@@ 3 Conclusion
By applying PIF to structure-based drug design, we develop MoIPIF which demonstrates superior performance in generating 3D molecules conditioned on protein binding pockets. Furthermore, we implement a geometry-enhanced training strategy that provides atomic context to assist model training process. Comprehensive evaluations demonstrate the framework’s effectiveness, showing significant improvements in both general molecular properties and local geometries compared to state-of-the-art baselines. Case studies on targets including 2v3r, 1l3l and HAT1 (6VO5) further validate MoIPIF’s ability to generate novel molecules with enhanced binding properties over reference compounds.
@@ 4.2 Parameter Interpolation Flow
3:$t\Leftarrow i/n$
4:$\boldsymbol{\theta}(\textbf{x}_{\text{data}})\in\{\textbf{a}\mid p(\textbf{x}|\textbf{a})=\delta(\textbf{x}-\textbf{x}_{\text{data}})\}$
5:$f(t)\Leftarrow 1-\gamma^{t}$
6:$\boldsymbol{\theta}_{t}\Leftarrow f(t)\boldsymbol{\theta}(\textbf{x}_{\text{data}})+(1-f(t))\boldsymbol{\theta}_{\text{prior}}$
@@
8:$\hat{\boldsymbol{\theta}}\Leftarrow\boldsymbol{\Phi}(\textbf{m})$
9:$\Delta t\Leftarrow 1/n$
10:$\boldsymbol{\theta}_{t+\Delta t}\Leftarrow f(t+\Delta t)\boldsymbol{\theta}(\textbf{x}_{\text{data}})+(1-f(t+\Delta t))\boldsymbol{\theta}_{\text{prior}}$
11:$\hat{\boldsymbol{\theta}}_{t+\Delta t}\Leftarrow f(t+\Delta t)\hat{\boldsymbol{\theta}}+(1-f(t+\Delta t))\boldsymbol{\theta}_{\text{prior}}$
@@
13:$\boldsymbol{\Phi}\Leftarrow\boldsymbol{\Phi}-\alpha\nabla_{\boldsymbol{\Phi}}L_{t}$
@@
2:$\boldsymbol{\theta}_{0}\Leftarrow\boldsymbol{\theta}_{\text{prior}}$

### v3

@@ 1 Introduction
Empirical evaluations conducted on the CrossDocked2020 dataset [36] demonstrate that MolPIF has comprehensive advantages across five key dimensions: (1) Superior de novo generation capability, producing molecules with superior binding properties and chemical validity; (2) Accurate geometric reproduction of molecular structural distributions, including rings, bond lengths, and bond angles; (3) Comprehensive chemical space modeling, with substantial coverage of 2D structural features and accurate 3D conformational distribution, while extending shape diversity beyond the reference distribution; (4) Flexible adaptation to different prior distributions, where we systematically compared and analyzed the advantages and limitations of using Gaussian versus Laplace distributions as priors for modeling atomic coordinates; (5) Effective lead optimization, demonstrating robust performance in enhancing drug candidate properties.
@@ 2.4.2 Analysis of MolPIF on local geometries
2.4.3 Analysis of MolPIF on chemical space distribution
In contrast to analyzing the local geometry of molecules generated by MolPIF, the examination of their chemical space distribution provided a more macroscopic perspective. Inspired by PMDM [55], we employed both 2D and 3D molecular descriptors—the Extended-Connectivity Fingerprints(ECFP), RDKit, and USRCAT ((Ultrafast Shape Recognition with CREDO Atom Types) [56]—to represent the chemical space of generated molecules and reference test set molecules, with particular emphasis on structural topology. Specifically, we adopted the ECFP, which implement the Morgan [57] algorithm to assign unique atom identifiers. These fingerprints encode atomic environments by integrating atom types (e.g., connectivity), chemical features (e.g., hydrogen bond donors/acceptors), and neighboring atomic contexts. The RDKit fingerprint, conceptually derived from the Daylight fingerprint, quantified 2D molecular substructures by evaluating atom and bond types. Conversely, USRCAT enhanced the Ultrafast Shape Recognition (USR) algorithm by integrating pharmacophoric descriptors to characterize 3D molecular shape.
The chemical space distribution visualized by t-SNE [58] was presented in Fig. 2a–c. While MolPIF-generated molecules basically covered the test set’s chemical space in 2D substructure representation (Fig. 2a,b), they showed superior density alignment in 3D conformational space (Fig. 2c). The generated molecules’ high-density regions precisely matched the test set’s concentrated areas in 3D space, demonstrating MolPIF’s ability to capture authentic conformational distributions. This accurate density reproduction in 3D space was particularly notable given molecular conformation complexity. The visualization confirmed that MolPIF maintained comprehensive 2D feature coverage while achieving physically meaningful 3D distributions that highlight the most relevant conformational regions observed experimentally.
Figure 2: Visualization of chemical space distribution and molecular shape characteristics. a-c, Chemical space distributions of molecules visualized via t-SNE in two-dimensional space, based on Morgan (a), RDKit (b), and USRCAT (c) fingerprints, respectively. d,e, Shape distributions are compared between generated molecules (d) and reference set molecules (e), represented using Normalized Principal Moment of Inertia ratios (NPR). f, Statistical comparison of Plane of Best Fit (PBF) descriptor values between generated molecules (n=10,000) and reference set (n=100), with box plots showing median (center line), interquartile range (box limits), 1.5×IQR whiskers, and extreme values.
To further analyze the 3D shape distribution of generated molecules, we employed molecular descriptors to characterize molecular structures beyond conventional fingerprints. Following the methodology used in PMDM, we utilized two widely recognized molecular descriptors: Principal Moments of Inertia (PMI) [59] and Plane of Best Fit (PBF) [60], which enable comprehensive investigation of molecular shapes. The PMI analysis quantitatively characterized molecular geometries by classifying them into rod-shaped, disc-shaped, or sphere-shaped configurations based on their inertial properties. Meanwhile, the PBF approach determined the optimal fitting plane through all heavy atoms in a given molecular conformation and subsequently computed the average deviation of these atoms from the reference plane, providing a complementary measure of molecular planarity.
Fig. 2d,e visualize the Normalized Principal Moment of Inertia ratios (NPR) on a ternary plot, where proximity to each corner reflects the dominance of rod-, disc-, or sphere-like morphologies. Notably, MolPIF-generated molecules exhibited a distribution closely aligned with the test set, with both populations predominantly clustered near the rod-like vertex. Importantly, MolPIF further extended the shape diversity beyond the reference’s coverage, as evidenced by generated samples reaching the disc- and sphere-like regions—areas sparsely represented (or absent) in the original test distribution. This demonstrated MolPIF’s dual capability: (1) faithfully learning the 3D molecular shape distribution of the datasets, and (2) exploring novel molecular structures. Moreover, as illustrated in Fig. 2f, the PBF value distribution of molecules generated by MolPIF showed agreement with that of the test set, demonstrating the model’s capability to accurately learn and reproduce the spatial planarity characteristics of authentic molecular conformations. This precise modeling of PBF metrics indicated that the model had successfully captured the intrinsic relationship between the spatial arrangement of heavy atoms and molecular planarity.
MolPIF effectively captured and extended the chemical space of reference molecules, achieving strong coverage in both 2D substructure and 3D conformational representations. While maintaining fidelity to dominant structural trends, it also explored novel shape diversity beyond the original distribution, as demonstrated by PMI and PBF analysis. This balance between accurate learning and innovative generation highlighted its potential for drug discovery applications requiring precise 3D molecular design.
@@ 2.4.4 Case analysis for the performance of MolPIF in de novo molecule generation
For the case study, we selected three representative binding pockets (2V3R, 1L3L, and 6VO5) for de novo molecule generation. Three-dimensional visualization of the generation results is presented in Fig. 3. Fig. 3a–d illustrate the generation outcomes for pockets 2V3R and 1L3L. The displayed molecules represented the top-performing candidates from 100 generated samples per model, selected according to a comprehensive ranking that considered all evaluation metrics. Quantitative analysis demonstrated that MolPIF-generated molecules exhibited superior performance across multiple metrics under the constraint of a fixed number of atoms. This improved ligand efficiency suggested substantial advantages for downstream optimization and experimental synthesis, substantiating the practical applicability and potential of MolPIF in real-world drug discovery applications.
To evaluate the model’s performance in a biologically relevant context, we applied MolPIF to generate potential inhibitors targeting HAT1. HAT1 is an enzyme that catalyzes the acetylation, which plays a critical role in various physiological processes, and its dysregulation has been closely associated with multiple human diseases, particularly cancers [61]. The generated molecules were then compared with both the reference molecule from the crystal structure and the preliminary screening hits reported in the literature. In this evaluation, MolPIF generated 100 candidate molecules targeting the HAT1 binding site (PDB ID: 6VO5). Remarkably, approximately 50% of these molecules outperformed molecule H9 in key metrics (Fig. 3e,f), with the vast majority showing improvement over the reference. Notably, H9 was identified from 100,000 PocketFlow-generated molecules [62] and has been experimentally validated for its biological activity ($IC_{\text{50}}=72.36\mu M$). These results indirectly demonstrated the exceptional efficiency of MolPIF in de novo generation of active molecules as starting points for novel targets.
@@ 2.4.6 The performance of MolPIF in lead optimization
MolPIF demonstrated applicability not only to de novo molecular generation tasks but also to atom-level lead optimization. This capability enabled arbitrary specification of fixed atoms within given molecules while allowing the model to generate desired substituent regions. Fig. 5 presents selected lead optimization cases performed on reference molecules targeting 1umd, 3ZCW, and 6KZZ.
The target 1umd was selected from the CrossDocked test set, with fixed ligand atom selection following the protocol established by CBGBench [54]. Fig. 5a displays some molecular structures generated by MolPIF for pocket 1umd during lead optimization. The reference ligand was partitioned into distinct regions for four subtasks: fragment growth, linker design, scaffold hopping, and side-chain decoration. MolPIF exhibited dual functional capabilities: (1) extending existing scaffold groups through Fragment and Side Chain modifications, and (2) integrating discrete fragments into complete molecular structures via Linker and Scaffold operations. Notably, the model accomplished molecular optimization without explicit gradient guidance from property prediction, relying solely on learned structural information. This suggested that the model implicitly captured structural features associated with favorable molecular properties. Such atom-level lead optimization facilitated straightforward modification or replacement of molecular substructures, enabling optimization of physicochemical properties or circumvention of patent restrictions, thereby demonstrating significant practical potential.
@@ 3 Conclusion
By applying PIF to structure-based drug design, we develop MoIPIF which demonstrates superior performance in generating 3D molecules conditioned on protein binding pockets. Furthermore, we implement a geometry-enhanced training strategy that provides atomic context to assist model training process. Comprehensive evaluations demonstrate the framework’s effectiveness, showing significant improvements in both general molecular properties and local geometries compared to state-of-the-art baselines. Chemical space analysis confirms MoIPIF’s ability to accurately replicate and expand molecular structural diversity. Case studies on targets including 2v3r, 1l3l and HAT1 (6VO5) further validate MoIPIF’s ability to generate novel molecules with enhanced binding properties over reference compounds.
@@ 4.2 Parameter Interpolation Flow
[…]
@@
[…]
@@
[…]